\begin{thebibliography}{0}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi
\expandafter\ifx\csname bibnamefont\endcsname\relax
  \def\bibnamefont#1{#1}\fi
\expandafter\ifx\csname bibfnamefont\endcsname\relax
  \def\bibfnamefont#1{#1}\fi
\expandafter\ifx\csname citenamefont\endcsname\relax
  \def\citenamefont#1{#1}\fi
\expandafter\ifx\csname url\endcsname\relax
  \def\url#1{\texttt{#1}}\fi
\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
\providecommand{\bibinfo}[2]{#2}
\providecommand{\eprint}[2][]{\url{#2}}

\end{thebibliography}


\begin{thebibliography}{99}
\bibitem{Biss} H. Biss, L. Sobirey, N. Luick, M. Bohlen, J.J. Kinnunen, G.M. Bruun, T. Lompe, H. Moritz, \href{https://doi.org/10.1103/PhysRevLett.128.100401}{``Excitation Spectrum and Superfluid Gap of an Ultracold Fermi Gas''}, Phys. Rev. Lett. {\bf 128}, 100401  (2022). 
\bibitem{LK} L. Landau, I. Khalatnikov, ``Teoriya vyazkosti Geliya-II'', Zh. Eksp. Teor. Fiz. {\bf 19}, 637 (1949).
\bibitem{Annalen} H. Kurkjian, Y. Castin, A. Sinatra, \href{https://doi.org/10.1002/andp.201600352}{``Three-phonon and four-phonon interaction processes in a pair-condensed Fermi gas''}, Annalen der Physik (Berlin) {\bf 529}, 1600352 (2017).
\bibitem{SonWingate} D.T. Son, M. Wingate, \href{https://doi.org/10.1016/j.aop.2005.11.001}{``General coordinate invariance and conformal invariance in nonrelativistic physics: Unitary Fermi gas''}, Ann. Physics (Amsterdam) {\bf 321}, 197 (2006).
\bibitem{Zwierlein} Mark J.H. Ku, A.T. Sommer, L.W. Cheuk, M.W. Zwierlein, \href{https://doi.org/10.1126/science.1214987}{``Revealing the Superfluid Lambda Transition in the Universal Thermodynamics of a Unitary Fermi Gas''}, Science {\bf 335}, 563 (2012).
\bibitem{RupakSchafer} G. Rupak, T. Schäfer, \href{https://doi.org/10.1016/j.nuclphysa.2008.11.004}{``Density functional theory for non-relativistic fermions in the unitarity limit''}, Nuclear Physics A {\bf 816}, 52 (2009).
\bibitem{Randeria} R.B. Diener, R. Sensarma, M. Randeria, \href{https://doi.org/10.1103/PhysRevA.77.023626}{``Quantum fluctuations in the superfluid state of the BCS-BEC crossover''}, Phys. Rev. A {\bf 77}, 023626 (2008).
\bibitem{concav} H. Kurkjian, Y. Castin, A. Sinatra, \href{https://doi.org/10.1103/PhysRevA.93.013623}{``Concavity of the collective excitation branch of a Fermi gas in the BEC-BCS crossover''}, Phys. Rev. A {\bf 93}, 013623 (2016).
\bibitem{Zwerger} R. Haussmann, M. Punk, W. Zwerger, \href{https://doi.org/10.1103/PhysRevA.80.063612}{``Spectral functions and rf response of ultracold fermionic atoms''}, Phys. Rev. A {\bf 80}, 063612 (2009).
\bibitem{CRAS} Y. Castin, A. Serafin, A. Sinatra \href{https://doi.org/10.5802/crphys.169}{``Amortissement des phonons dans un superfluide 2D : insuffisance de la règle d’or de Fermi à basse température''}, Comptes Rendus Physique {\bf 24}, 187 (2023).
\bibitem{PRL} Y. Castin, A. Sinatra, H. Kurkjian, \href{https://doi.org/10.1103/PhysRevLett.119.260402}{``Landau phonon-roton theory revisited for superfluid ${}^4$He and Fermi gases''}, Phys. Rev. Lett. {\bf 119}, 260402 (2017).
\bibitem{NishidaSon}  Y. Nishida, D.T. Son, \href{https://doi.org/10.1103/PhysRevLett.97.050403}{``$\epsilon$ Expansion for a Fermi gas at infinite scattering length''}, Phys. Rev. Lett. {\bf 97}, 050403 (2006).
\bibitem{Ketterle} A. Schirotzek, Y.I. Shin, C.H. Schunck, W. Ketterle, \href{https://doi.org/10.1103/PhysRevLett.101.140403}{``Determination of the superfluid gap in atomic Fermi gases by quasiparticle spectroscopy''}, Phys. Rev. Lett. {\bf 101}, 140403 (2008).
\end{thebibliography}

\documentclass[twocolumn,floatfix,aps,prl]{revtex4}
\usepackage{hyperref}
\hypersetup{breaklinks=true}
\hypersetup{colorlinks=true}
\hypersetup{urlcolor=blue}
\hypersetup{citecolor=blue}
\hypersetup{linkcolor=blue}
\begin{document}
\title{Comment on ``Excitation Spectrum and Superfluid Gap of an Ultracold Fermi Gas''}
\author{Yvan Castin}
\affiliation{Laboratoire Kastler Brossel, ENS-Universit\'e PSL, CNRS, Universit\'e Sorbonne and Coll\`ege de France, 24 rue Lhomond, 75231 Paris, France}
\maketitle

At zero temperature, as any three-dimensional superfluid with short-range interactions, a gas of paired fermionic atoms exhibits an acoustic excitation branch of low-wavenumber expansion $\omega_q=c q [1+\zeta q^2/k_{\rm F}^2+O(q^4\ln q)]$ with $c$ the speed of sound and $\zeta$ the curvature parameter (scaled by the Fermi wavenumber $k_{\rm F}$). Reference \cite{Biss} claims to have experimentally determined whether the branch has a convex $\zeta>0$ or concave $\zeta<0$ start, depending on the interaction strength. This is crucial information that dictates the nature of the gas relaxation mechanisms at low temperatures (the well studied three-phonon Beliaev-Landau or the yet unobserved four-phonon Landau-Khalatnikov mechanism \cite{LK,Annalen}). However, in this Comment, we argue that the high wavenumbers $q$ and temperature $T$ used in the experiment introduce a large bias capable of turning a convex branch into a concave one, so that further measurements at lower $q$ and $T$ are required to give a definitive answer.

To fix ideas, we consider the unitary limit $1/k_{\rm F}a=0$ with $a$ the $s$-wave scattering length, where the interactions are strongest and no solid argument can predict the sign of $\zeta$. Theoretically, one has $\zeta=-\pi^2$ $(2\xi_{\rm B})^{1/2}$ $[c_1+(3/2) c_2]$ where Bertsch's parameter gives the chemical potential in units of the Fermi energy $\mu=\xi_{\rm B} E_{\rm F}$ and $c_{1,2}$ quantify gradient corrections to quantum hydrodynamics \cite{SonWingate}. Only $\xi_{\rm B}$ is well known, $\xi_{\rm B}\simeq 3/8$ \cite{Zwierlein}. The dimensional expansion in powers of $\epsilon=4-d=1$ gives $c_1\simeq -0.0624(1-2\epsilon/3)+O(\epsilon^2)$ and $c_2=O(\epsilon^2)$ \cite{RupakSchafer}, so $\zeta>0$ to subleading order. Anderson's RPA, spectrally equivalent to the Gaussian fluctuations approximation of \cite{Randeria}, also predicts a positive value $\zeta_{\rm RPA}\simeq0.0838$ \cite{concav} (for $c_1\simeq-0.021$ \cite{RupakSchafer} this gives $c_2\simeq 0.0073\ll|c_1|$). The experimental value $\zeta_{\rm exp}=-0.085(8)$ \cite{Biss} is negative. However, assuming that the RPA is correct and that the branch start is convex, as we will do, actually has no clear incompatibility with the experiment, because the analysis in \cite{Biss} suffers from two serious limitations.

First, the value $\zeta_{\rm exp}$, obtained by cubic fitting of $\omega_q$ \cite{Biss}, could strongly depend on the fitting interval if too wide. In the RPA, fitting $\omega_q$ e.g. to the interval $0.22\leq q/k_{\rm F}\leq 1.08$ of Figure 1 in \cite{Biss} gives $\zeta_{\rm RPA}^{\rm fit}\simeq -0.026$, which even has the wrong sign. Since $\omega_q^{\rm RPA}$ has an inflection point at $q/k_{\rm F}\simeq 0.5$ \cite{concav} the fit blindly mixes convex and concave parts, which also explains the erroneous (negative) value of $\zeta_{\rm RPA}$ in \cite{Zwerger}.

Second, The high temperature $T\simeq 0.13 T_{\rm F}\simeq 0.8 T_{\rm c}$ ($T_{\rm c}$ is the superfluid transition temperature) in \cite{Biss} could modify the curvature of the acoustic branch by a non-negligible amount $\delta\zeta^{\phi\phi}$ via interaction with thermal phonons $\phi$. Treating the cubic phonon-phonon coupling $H_3^{\phi\phi}$ to second order and the quartic coupling $H_4^{\phi\phi}$ to first order, then taking the limit $k_{\rm B} T/mc^2\to 0$ ($m$ is the fermion mass), \cite{CRAS} obtains an expression for the thermal shift of $\omega_q$. This gives $\delta\zeta^{\phi\phi}\sim-[\pi^2/(3\xi_{\rm B})^{3/2}](T/T_{\rm F})^2$, or $\delta\zeta^{\phi\phi}\simeq -0.140$ at the experimental temperature. Since the small parameter used $k_{\rm B}T/mc^2\simeq 0.5$ is not $\ll 1$, we abandon the $T\to 0$ limit and add corrective curvature factors $(1\pm\alpha q^2/k_{\rm F}^2)$ to the amplitudes $\rho_q$ and $\phi_q$ of the superfluid density and phase quantum fluctuations ($\alpha=\pi^2(\xi_{\rm B}/2)^{1/2}[c_1-(3/2)c_2]\simeq-0.136$ \cite{Annalen}). We find $\delta\zeta^{\phi\phi}\simeq -0.110$, still negative enough to change the sign of curvature in the RPA. Furthermore, thermal pair dissociation creates fermionic quasiparticles $\gamma$ (another gas excitation branch) that interact with the phonons. Treating to second order the coupling $H_3^{\phi\gamma}$ and to first order the coupling $H_4^{\phi\gamma}$ given in \cite{PRL} with curvature factors in $\omega_q$, $\rho_q$ and $\phi_q$, we find $\delta\zeta^{\phi\gamma}\simeq -0.052$. This is a rough estimate: \cite{PRL} uses a simple local-density approximation whose small parameter $(k_{\rm B} T/m_*c^2)^{1/2}$ is $\approx1$ here (we use the $\gamma$ dispersion relation of \cite{NishidaSon} with the effective mass $m_*\simeq 0.56m$ and an energy minimum $\Delta=0.44 E_{\rm F}$ located at wavenumber $k_0=0.92 k_{\rm F}$ \cite{Ketterle}). Summing the thermal corrections gives $\zeta_{\rm RPA}^{\rm th}\simeq -0.078$ to compare with $\zeta_{\rm exp}=-0.085(8)$. 

The experiment, at first sight at odds with the RPA, could therefore very well be in agreement with it due to large finite-momentum and thermal bias, and the sign of curvature announced in \cite{Biss} may differ from the zero-temperature low-momentum one relevant for phonon damping.

{\sl Acknowledgments} -- We thank Alice Sinatra for her comments on the text.

\end{document}